A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
Answer by ankor@dixie-net.com(22740)

A coffee merchant has coffee beans that sell for $9 per pound and $12 per pound. The two types are to be mixed to create 100lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
:
Let x = amt of $12 coffee required
It said that a total of 100lb is required, therefore
Let (100-x) = amt of $9 coffee required
:
12x + 9(100-x) = 11.25(100)
12x + 900x - 9x = 1125
12x - 9x = 1125 - 900
3x = 225
x = 225/3
x = 75 lb of $12 coffee, obviously that leaves 25 lb of $9 coffee required
:
:
Check solution in original equation
12(75) + 9(25) = 1125
:
:
Not that hard, right?